OpenStudy (anonymous):

Write the complex number in the form a + bi. 3(cos 270° + i sin 270°) Could someone please explain how to do problems like these?

OpenStudy (uri):

Just evaluate cos(270) and sin(270) and do any simplifying.

OpenStudy (anonymous):

3(0+-1i)=0-3i=-3i is this what you looking for?

OpenStudy (anonymous):

@burn23 yes!

OpenStudy (mathmath333):

\(\large \tt \begin{align} \color{black}{z=a+bi\\~\\ r=|z|=\sqrt{a^2+b^2}\\~\\ \alpha=tan^{-1}(\dfrac{b}{a})\\~\\ trig form \\~\\ z=r(cos\alpha+isin\alpha)\\~\\ } \end{align}\)

OpenStudy (mathmath333):

go reverse

OpenStudy (anonymous):

you gone give me a medal?:)